### v2

# Argumentatively Coherent Judgmental Forecasting

Abstract
Judgmental forecasting employs human opinions to make predictions about future events, rather than exclusively historical data as in quantitative forecasting. When these opinions form an argumentative structure around forecasts, it is useful to study the properties of the forecasts from an argumentative perspective. In this paper, we advocate and formally define a property of argumentative coherence, which, in essence, requires that a forecaster’s reasoning is coherent with their forecast. We then conduct three evaluations with our notion of coherence. First, we assess the impact of enforcing coherence on human forecasters as well as on Large Language Model (LLM)-based forecasters, given that they have recently shown to be competitive with human forecasters. In both cases, we show that filtering out incoherent predictions improves forecasting accuracy consistently, supporting the practical value of coherence in both human and LLM-based forecasting. Then, via crowd-sourced user experiments, we show that, despite its apparent intuitiveness and usefulness, users do not generally align with this coherence property. This points to the need to integrate, within argumentation-based judgmental forecasting, mechanisms to filter out incoherent opinions before obtaining group forecasting predictions.
1 Introduction
Judgmental forecasting involves the use of human opinion to predict future events, as opposed to relying solely on historical data as in quantitative forecasting Lawrence et al. (2006); Zellner et al. (2021). In the area of judgmental forecasting, Tetlock et al. (2014) established techniques for supporting forecasting, and Tetlock and Gardner (2016) improved these techniques by selecting “Superforecasters” (forecasters who naturally excel at making predictions). However, judgmental forecasting can be influenced by irrationality and cognitive biases Kahneman (2011), which can lead to inaccurate predictions. The importance of coherence in this context cannot be overstated Karvetski et al. (2022). For instance, if a forecaster allows their own personal biases to influence their prediction, it could lead to skewed, and thus less accurate, results.
To overcome irrationality and biases, many solutions have been proposed over the years, e.g. Wallsten and Budescu (1995); Arkes (2001); Chang et al. (2016); Zellner et al. (2021), including some solutions Karvetski et al. (2013); Fan et al. (2019); Gibson et al. (2021); Thomson et al. (2019); Ho et al. (2024) showing that imposing coherence improves forecasting accuracy. Further, coherence constraints has been shown to help in hierarchical forecasting Rangapuram et al. (2021); Panagiotelis et al. (2023); Rangapuram et al. (2023); Olivares et al. (2023). Differently from these approaches, in this paper we propose a novel solution based on a notion of coherence deployed alongside debates (on forecasting questions) structured as in computational argumentation (see Atkinson et al. (2017); Baroni et al. (2018) for overviews). This field of research has emerged in recent years in AI as a methodology to conjugate representational needs of user-related cognitive models with computational models for automated reasoning Lippi and Torroni (2016). These computational models are formalised as argumentation frameworks, which excel in reasoning with incomplete information and for resolving conflicts. Within this landscape, existing research Irwin et al. (2022); Gorur et al. (2023); Toni et al. (2023) attempts to address the aforementioned issues with judgmental forecasting with the use of argumentation frameworks. Some of these works Toni et al. (2023) do not incorporate any coherence constraints. Other work Irwin et al. (2022), while showing promising results with the introduction of coherence constraints, focus on restricted settings, notably they can only handle binary questions and disallow the same arguments to be used for different questions, which does not maximise the potential of argumentation. Further work Gorur et al. (2023) addresses these limitations while imposing coherence constraints, but shows limitations as regards structuring and evaluating forecasters’ judgments (for example, forecasters may express no opinion but their judgment may be different from neutral).
In this paper, we make several contributions.
  • We define a novel, (arguably) natural notion of coherent forecaster behaviour, given predictions and votes on arguments, for a variant of the forecasting method of Gorur et al. (2023) addressing its limitations as regards structuring and evaluating forecasters’ judgments. Differently from existing notions, our notion of coherence is adaptable to different forecasting contexts through controllable parameters.
  • Then, to assess the usefulness of our notion of coherence in practice, we carry out three evaluations.
    1. 1.
      We conduct an in-person user study with a variant of the argumentation-based forecasting system ArguCast111argucast.herokuapp.com (this variant implements our variant of the forecasting method of Gorur et al. (2023)). With this study, we aim to see whether applying coherence gets predictions closer to state-of-the-art forecasting systems.
    2. 2.
      We perform experiments on Large Language Model (LLM)-based forecasters, given that they have recently shown to be competitive with human forecasters Halawi et al. (2024). We use Argumentative LLMs (ArgLLMs) Freedman et al. (2025) to first generate arguments for and against the forecasting arguments, and then we apply coherence to see whether forecasting accuracy improves.
    3. 3.
      We run a crowd-sourced user study where we assign users variants of debates, formulated with our forecasting method, and assess their alignment with our notion of coherence.
Results from the in-person user study and from the experiments with the LLMs show that applying coherence provides consistent improvements on forecasting accuracy, thus showing the practical need for coherence constraints. Results from the crowd-sourced study demonstrate that users only partially align with our notion. Thus, given the benefits of coherence towards accuracy, this final study points to the need to enforce coherence constraints for filtering out incoherent opinions in ArguCast, and potentially also in other argumentation-based forecasting tools (e.g. Irwin et al. (2022)).
This pre-print is an extended version of Gorur et al. (2025), including supplementary material.
2 Related Work
Judgmental Forecasting Baselines
We measure our method, filtering out the opinion of incoherent forecasters, against a baselines representing the state-of-the-art in (judgmental) forecasting. These are Metaculus222www.metaculus.com and prediction markets (in particular we use Betfair333www.betfair.com).
Coherence in Judgmental Forecasting
In Judgmental forecasting, there is some work that tries to incorporate coherence, for example by investigating the best ways of eliciting probabilities from humans Wallsten and Budescu (1995), how incentives and training change forecasters’ abilities Chang et al. (2016), the effect of scoring rules on forecasters Zellner et al. (2021), and how principles help overcome overconfidence Arkes (2001). Some existing solutions Karvetski et al. (2013); Fan et al. (2019); Gibson et al. (2021) show that forecasting accuracy improves by weighting judges’ assessments based on their degree of coherence. Further, Thomson et al. (2019) develop a measure of coherence to identify the degree of independence between sets of forecasts, thus improving the aggregated forecasts. Additionally, Ho et al. (2024) introduce a notion of probabilistic coherence that identifies superior forecasters. Finally, there has been some work on coherence constraints in hierarchical forecasting Rangapuram et al. (2021); Panagiotelis et al. (2023); Rangapuram et al. (2023); Olivares et al. (2023). In this paper we do not compare our notion against existing approaches, since they do not operate in our setting.
Forecasting and Argumentation
Existing research Irwin et al. (2022); Gorur et al. (2023); Toni et al. (2023) explores the integration of computational argumentation for eliciting, representing and evaluating the argumentative reasoning of users. Specifically, Irwin et al. (2022) were the first to apply computational argumentation to support judgmental forecasting. They introduced their notion of coherence based on how forecasters update their predictions and their reasoning behind those updates. However, as mentioned in the introduction, their underlying framework exhibits some limitations in that it can only handle binary questions and does not allow the same arguments to be used for different questions. It is also important to note that our notion of coherence is based solely on forecasters’ argumentative reasoning and their predictions rather than prediction updates. Further demonstrating the benefits of argumentation in this context is Toni et al. (2023), which incorporates users’ opinions and probabilistic estimates to aid forecasting, but without considering coherence constraints. In addition, the role of argumentation for forecasting is indirectly advocated by resources such as the FORECAST2023 dataset Gorska et al. (2024), where the rationales from the Hybrid-Forecasting Competition are annotated with support/attack relations.
We use a variant of ArguCast Gorur et al. (2023), which addressed the shortcomings of Irwin et al. (2022). Our variant exhibits additional properties (e.g. neutral and bias-based forecasting) for the underlying formalism for structuring and evaluating forecasters’ judgments (see Appendix A). Also, our notion of coherence is parameterised and adaptable to different forecasting contexts, whereas the notion of coherence in Gorur et al. (2023) is fixed.
LLMs and Forecasting
Recent research demonstrates the potential of LLMs in forecasting, e.g. in time-series forecasting, LLMs have been shown to improve accuracy Jin et al. (2024); Tang et al. (2024); Liu et al. (2024). However, direct use of LLMs in judgmental forecasting has yielded mixed results: while some studies Schoenegger and Park (2023); Abolghasemi et al. (2023) show that using these models directly does not improve forecasting accuracy, others Halawi et al. (2024) have concluded that the current LLMs are nearly as good as human forecasters if they are supported with Retrieval Augmented Generation (RAG) Zhao et al. (2024). They further point that the aggregation of forecasts from humans and LLMs with RAG achieve the best forecasting accuracy. Combining human insights with LLM-supported advice has also proven effective in accuracy Schoenegger et al. (2024). Our approach in Section 6 incorporates LLMs on forecaster data to improve forecasting accuracy. It does so by adapting, towards forecasting, ArgLLMs Freedman et al. (2025), an existing system for mining argumentation frameworks from LLMs to verify claims.
3 Background
In this section, we give the relevant background on computational argumentation, including for forecasting.
A Quantitative Bipolar Argumentation Framework (QBAF) Baroni et al. (2019) is a tuple $\langle\mathcal{X},\mathcal{A},\mathcal{S},\tau\rangle$ where $\mathcal{X}$ is a set of arguments; $\mathcal{A}\!\subseteq\!\mathcal{X}\times\mathcal{X}$ is a relation of attack; $\mathcal{S}\!\subseteq\!\mathcal{X}\times\mathcal{X}$ is a relation of support; and $\tau\!:\!\mathcal{X}\!\rightarrow\![0,1]$ is a total function, with $\tau(a)$ the base score of $a\!\in\!\mathcal{X}$.
Example 1.
Figure 1 depicts as a graph the QBAF $\langle\mathcal{X},\mathcal{A},\mathcal{S},\tau\rangle$ where $\mathcal{X}=\{a,b,c\}$, $\mathcal{S}=\{(a,b)\}$, and $\mathcal{A}=\{(a,c)\}$ with $\tau(a)=0.5$, $\tau(b)=0.1$, and $\tau(c)=0.7$.
Figure 1: The QBAF in Example 1 as a graph, with arguments represented as nodes, supports as green solid edges, and attacks as red dashed edges.
QBAFs are equipped with so-called gradual semantics, assessing the (dialectical) strength of arguments. We use the Discontinuity-Free Quantitative Argumentation Debate (DF-QuAD) gradual semantics Rago et al. (2016) to obtain the strength of the forecasters’ opinions. For a given QBAF $\langle\mathcal{X},\mathcal{A},\mathcal{S},\tau\rangle$, for any $x\in\mathcal{X}$ with $n\geq 0$ attackers with strengths $v_{1},\ldots,v_{n}$, $m\geq 0$ supporters with strengths $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ and $\tau(x)=v_{0}$, DF-QuAD computes $x$’s strength as $\sigma(x)\!=\!\mathcal{C}(v_{0},\mathcal{F}(v_{1},\ldots,v_{n})$, $\mathcal{F}(v_{1}^{\prime},\ldots,v_{m}^{\prime}))$, where, for any $w_{1},\ldots,w_{k}$, $\mathcal{F}(w_{1},\ldots,w_{k})$ is 0 if k=0k\!=\!0 and $1-\prod_{i=1}^{k}(|1-w_{i}|)$ otherwise and $\mathcal{C}$ is defined as follows: for $v_{a}=\mathcal{F}(v_{1},\ldots,v_{n})$ and $v_{s}=\mathcal{F}(v_{1}^{\prime},\ldots,v_{m}^{\prime})$, if $v_{a}=v_{s}$ then $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}$; else if $v_{a}>v_{s}$ then $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}-(v_{0}\cdot|v_{s}-v_{a}|)$; otherwise $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}+((1-v_{0})\cdot|v_{s}-v_{a}|)$.
Example 2.
Consider the QBAF from Example 1. Using DF-QuAD, $\sigma(b)=\tau(b)=0.1$ and $\sigma(c)=\tau(c)=0.7$ given that $b,c$ are neither attacked nor supported. Then, as $b$ is the single supporter and $c$ is the single attacker of $a$, and the latter is stronger than the former, $\sigma(a)=\mathcal{C}(0.5,0.7,0.1)=0.5-(0.5\cdot 0.6)=0.2$.
[…]
Example 3.
Consider an ACF $\langle\mathcal{X,R,U,V,P}\rangle$ with the same arguments and relations as the QBAF from Example 1, $\mathcal{U}=\{u\}$ with a single user disagreeing with $b$ (i.e., $\mathcal{V}(u,b)=$-) and agreeing with $c$ (i.e., $\mathcal{V}(u,c)=+$), and user $u$’s prediction on the forecasting argument given by $\mathcal{P}(u,a)=0.85$. This is visualised in Figure 2.
Figure 2: The ACF from Example 3 as a graph, with the forecasting argument represented as a cyan outlined node, the other arguments as gray outlined nodes, supports as green solid edges and attacks as red dashed edges.
For the remainder of the paper, unless specified otherwise, we assume as given an ACF $\langle\mathcal{X,R,U,V,P}\rangle$, with components as above.
4 Argumentative Coherence of Forecasters
In this section, we define a novel way of identifying incoherent forecasters, based on their argumentative opinions and predictions. First, we define a novel notion of forecasters as QBAFs drawn from ACFs:
Definition 1.
A forecaster QBAF for $u\in\mathcal{U}$ is a QBAF $\langle\mathcal{X},\mathcal{A}_{u},\mathcal{S}_{u},\tau_{u}\rangle$ such that, for $a,b,c\in\mathcal{X}$:
$\begin{split}\mathcal{S}_{u}=\{(b,c)\in\mathcal{R}|((b,c)\in\mathcal{S}\land(\mathcal{V}(u,b)=&\mathcal{V}(u,c)\\ \lor(\mathcal{V}(u,b)\neq-\land&\mathcal{V}(u,c)\neq-))\\ \lor((b,c)\in\mathcal{A}\land(\mathcal{V}(u,b)=-\land&\mathcal{V}(u,c)\neq-)))\}\end{split}$
$\begin{split}\mathcal{A}_{u}=\{(b,c)\in\mathcal{R}|((b,c)\in\mathcal{A}\land(\mathcal{V}(u,b)=&\mathcal{V}(u,c)\\ \lor(\mathcal{V}(u,b)\neq-\land&\mathcal{V}(u,c)\neq-))\\ \lor((b,c)\in\mathcal{S}\land(\mathcal{V}(u,b)=-\land&\mathcal{V}(u,c)\neq-)))\}\end{split}$
$\tau_{u}(a)\begin{cases}\in[0,1]&\text{if}\ a\in\mathcal{F}\\ =0.5&\text{if}\ \mathcal{V}(u,a)\in\{+,-\},\\ =0&\text{if}\ \mathcal{V}(u,a)\in\{?,\text{undefined}\}.\end{cases}$
In a forecaster QBAF, we define $\mathcal{A}_{u}$ and $\mathcal{S}_{u}$ based on attack and support relations in ACFs: if a forecaster disagrees with an argument, we assume their opinion is the negation of that argument; thus, the relations to and from that argument are flipped, leading the argument to have the opposite effect than originally considered. We keep the stance of a (child, parent) relation in $\mathcal{R}$ the same if the user has the same vote on both the child and parent argument or both of the arguments are not disagreed with. We change the stance of the relation if either one of the child or parent arguments is disagreed with and the other one is agreed with or is undecided (? or undefined).
The base score for the forecasting argument represents the confidence level of the forecaster on this argument prior to debating/voting. We choose the [0,1] scale for base scores in-line with existing work in gradual semantics Baroni et al. (2019). Higher assignments of this score mean there is stronger credibility for the forecasting argument, while lower assignments mean the forecasting argument is treated with more skepticism, and intermediate values provide a gradual scale between the two. Varying this base score allows us to control how much weight is given to the forecasting argument independently of the influence of supports and attacks. Also, in a forecaster QBAF, an argument’s base score is 0.5 when the forecaster agrees or disagrees with the argument (we use 0.5 as it is the middle point between 0 and 1). Finally, in a forecaster QBAF, we set the base score to 0 for an argument that the user has not voted on, as we do not want the argument with no opinion to have an effect on the forecasting argument’s strength. The following example illustrates considerations for how the base score for forecasting arguments can be assigned.
Example 4.
Consider a forecasting argument ff="X will win their next professional tennis match". We may assume a neutral assignment to the forecasting argument (e.g. $\tau(f)=0.5$) as we cannot make any assumptions on the initial confidence. If we know X is an amateur tennis player, then we would expect a lower confidence level, hence a lower assignment to the forecasting argument (e.g. $\tau(f)=0.03$). If instead we know X is the top-ranked player this season, then we would expect a higher confidence level, hence a higher assignment to the forecasting argument (e.g. $\tau(f)=0.7$). Thus, the base score reflects facts known to the forecaster prior to debate.
Our forecaster QBAFs differ from those underpinning the original ArguCast Gorur et al. (2023), specifically by satisfying properties of neutral forecasting (ensuring unbiased defaults when no opinions are expressed) and bias-based forecasting (ensuring opinions properly influence prediction strength), which Gorur et al. (2023) fails to satisfy (see Appendix A).
In the remainder, unless specified otherwise, we assume as given a forecaster QBAF $\langle\mathcal{X},\mathcal{A}_{u},\mathcal{S}_{u},\tau_{u}\rangle$ for $u\in\mathcal{U}$ in the given ACF.
We introduce a notion to determine whether a forecaster is argumentatively coherent. This notion constrains a forecasters’ prediction to align with their internal reasoning. In particular, we assess the consistency between the forecasters’ opinion (represented by the final strength of the forecasting argument within their QBAF), and their prediction, given two threshold values. This is with the intention of filtering out users who do not satisfy this condition, to potentially gain more accurate forecasts.
Definition 2.
Given functions $\xi_{1}:\mathcal{F}\rightarrow(0,1)$ and $\xi_{2}:\mathcal{F}\rightarrow(0,1)$ [^4] We disallow threshold values of exactly 0 or 1 as then the coherence condition would be too brittle in practice, since any minimal deviation would violate coherence., forecaster $u$ is prediction coherent iff $\forall{f\in\mathcal{F}}$:
  • if $\sigma(f)<\xi_{1}(f)$, then $\mathcal{P}(u,f)<\xi_{2}(f)$;
  • if $\sigma(f)>\xi_{1}(f)$, then $\mathcal{P}(u,f)>\xi_{2}(f)$;
  • if $\sigma(f)=\xi_{1}(f)$, then $\mathcal{P}(u,f)\in[\xi_{2}(f)-\epsilon,\xi_{2}(f)+\epsilon]$ for some small $\epsilon$.
This notion is more general than the one in Gorur et al. (2023) in that rather than fixing coherence thresholds, we introduce two threshold functions $\xi_{1}$ and $\xi_{2}$, providing greater flexibility across forecasting contexts.
In the remainder, unless specified otherwise, $f$ will indicate a generic forecasting argument.
Intuitively, $\xi_{1}$ is a function which sets the midpoint of the strength of an argument which arbitrates whether the reasoning attacking or supporting it is more dominant. In the experiments, with arguments’ base scores set up as in Definition 2, we use $\xi_{1}(f)=0.5$ throughout and leave to future work the study of other options. Meanwhile, $\xi_{2}$ is a function that gives the initial or expected likelihood of the forecasting argument occurring, e.g. as in Benjamin et al. (2023); Tetlock et al. (2014). In simple settings or when no prior knowledge is available, $\xi_{2}(f)$ may be set to 0.5, representing complete uncertainty, whereas higher or lower values may be set to the initial/expected probability, e.g. derived from expert judgment or prediction markets. Thus, $\xi_{2}(f)$ ensures that the forecasters’ reasoning is grounded in an appropriate starting point.
Example 5.
Consider the forecasting argument from Example 4, where the expected probability is a 50% chance of X winning their next match, and thus we would assign a neutral value ($\xi_{2}(f)=0.5$). Now consider another forecasting argument ‘X will win the professional tennis competition’, the probability is relatively low due to many competing players, so we may assign a lower value (e.g. $\xi_{2}(f)=0.2$). Finally, suppose we have a third forecasting argument ‘X will win a set in their next match’, the expected probability is higher, and thus we would assign a higher value ($\xi_{2}(f)=0.8$).
Both threshold values are task-specific and thus need to be optimised for each task. In the experiments, we use $\xi_{2}(f)=0.5$ or the initial probabilities from other judgmental forecasting platforms.
Finally, according to Definition 2, a forecaster is considered coherent if their prediction aligns with their opinion on the forecasting argument. So, if they mostly support (or attack) the forecasting argument then their prediction should be higher (or lower) than the initial or expected probability.
5 In-Person Experiments
In this section we detail the findings from our first user study, which we undertook with members of the public attending the Great Exhibition Road Festival (GERF) in London in June 2024555www.greatexhibitionroadfestival.co.uk. The participants used the ArguCast system with pre-prepared questions on sports, politics and finance. To assess the usefulness of coherence, we then compared the outcomes of our variant of ArguCast (from now on referred to as ArguCast for simplicity) with those of baseline systems (seen as the ground truth here, considering the much larger basis of forecasters therein, compared to ArguCast).
5.1 Research Question
We evaluated the following Research Question (RQ):
RQ1 Do predictions from ArguCast align more with the predictions from other systems when coherence is applied?
We expected the null hypothesis (i.e., a negative answer to the question) to be rejected.
5.2 Experimental Design
To explore RQ1, we asked participants to generate debates, express their opinions on the debates, and make predictions (in ArguCast) about three future events, along the following lines: ‘Will England win Euro 2024?’ (Euro), ‘Will the Labour party get a 50% majority on the 4th of July 2024 UK general election?’ (Labour), and ‘Will Nvidia become the most profitable company in the world before the end of September 2024?’ (Nvidia). Participants were given electronic devices such as tablets, laptops, and displays to use ArguCast. The generated debates are illustrated in Appendix B, and the forecaster QBAFs and predictions are available in the supplementary material. Then, we applied our notion of coherence with $\xi_{1}(f)=0.5$ and two options for $\xi_{2}$: $\xi_{2}(f)=0.5$, which is a neutral assignment, and $\xi_{2}(f)=\gamma$, where $\gamma$ is the initial prediction, which we obtained from prediction markets and Metaculus [^6] Specifically, the initial predictions for the Euro and Labour questions were taken from implied probabilities given by www.betfair.com, and the baseline for the Nvidia question was taken from the prediction for this event on www.metaculus.com on the 19th of June 2024.. We then computed the simple average of both the raw predictions (made by users) and coherent predictions (the predictions not filtered-out by coherence), and compared these predictions with the initial predictions by the baselines to assess the effect of the different values of $\xi_{2}$.
5.3 Participant Recruitment and Ethics Statement
We recruited 169 anonymous English-speaking participants at the GERF. [^7] Some participants focused only on some of the questions, see Table 1. The data collection was anonymous as we created anonymous accounts for the participants on ArguCast. After reading the participant information sheet, participants who used the ArguCast system gave implied consent. Ethics approval was obtained from our institution prior to conducting the experiments.
5.4 Results
The results are shown in Table 1.888Due to the small sample size and exploratory design, we did not perform statistical testing. For the Euro question, the raw average prediction (38.06) was far from the initial prediction (20). Applying coherence with $\xi_{2}(f)\!=\!0.5$ slightly moved the prediction further from the initial prediction, increasing it to (39.03). In contrast, coherence with $\xi_{2}(f)\!=\!\gamma$ brought the prediction down to 34.89, closer to the initial prediction. These results were achieved by filtering out 52 and 56 participants for $\xi_{2}(f)\!=\!0.5$ and $\xi_{2}(f)\!=\!\gamma$, respectively. For the Labour question, we see a similar trend where the raw average prediction (60.69) was far from the initial prediction (95). Applying coherence improved the results, giving 61.40 with $\xi_{2}(f)\!=\!0.5$ and 63.33 with $\xi_{2}(f)\!=\!\gamma$. It should be noted that achieving these corrections required filtering out more than half of the participants: out of 54, 29 for $\xi_{2}(f)\!=\!0.5$ and 48 for $\xi_{2}(f)\!=\!\gamma$, indicating that a substantial proportion of incoherent responses diverged from the initial prediction. For the Nvidia question, the raw average prediction (59.84) was lower than the initial prediction (80). Applying coherence improved on the raw average predictions, 65.10 with $\xi_{2}(f)\!=\!0.5$ and 77.00 with $\xi_{2}(f)\!=\!\gamma$. Again we filtered out more than half of the participants, 18 for $\xi_{2}(f)\!=\!0.5$ and 6 for $\xi_{2}(f)\!=\!\gamma$.
Initial ($\gamma$) Raw Avg. Coherent
$\xi_{2}(f)=0.5$ $\xi_{2}(f)=\gamma$
Euro 20 38.06 (N=83) 39.03 (N=31) 34.89 (N=27)
Labour 95 60.69 (N=54) 61.40 (N=25) 63.33 (N=6)
Nvidia 80 59.84 (N=38) 65.10 (N=20) 77.00 (N=6)
Table 1: Predictions for the three forecasting questions comparing the initial prediction (by prediction markets or Metaculus), with the raw average of participant predictions on ArguCast, and the coherent predictions with $\xi_{2}(f)\!=\!0.5$ or $\xi_{2}(f)\!=\!\gamma$. NN is the number of participating users.
These results indicate that filtering out incoherent predictions can improve aggregate forecasts. Overall, across all three questions, applying coherence with $\xi_{2}(f)=\gamma$ to ArguCast consistently moved predictions closer to the initial predictions. However, when coherence with $\xi_{2}(f)=0.5$ was applied the prediction moved further from the initial prediction for the Euro question. This suggests that using a neutral coherence threshold can still allow some of the incorrect predictions to be seen as coherent, thus pointing to the importance of a customisable notion of coherence. The results for the Labour and Nvidia questions further support this interpretation, as applying coherence with $\xi_{2}(f)=\gamma$ led to greater improvements over compared to applying rationality with $\xi_{2}(f)=0.5$. The consistency in these improvements is valuable, particularly given the complexity and noise inherent in human judgment data Kahneman (2011). Overall, we can reject the null hypothesis for RQ1, since applying coherence (for appropriate choices of the parameters improves forecasting accuracy.
6 LLM Experiments
In this section, we present the findings of an empirical experiment conducted using LLMs on forecasting questions from GJOpen (a public judgmental forecasting platform 999www.gjopen.com).
6.1 Research Question
Through these experiments, we address the following RQ:
RQ2 Does accuracy improve when coherence is applied?
We expected the null hypothesis (i.e., a negative answer to the question) to be rejected.
6.2 Experimental Setup
6.2.1 Dataset
We scraped 595 questions from GJOpen. Some of these questions included both binary and multiple-choice answers, yielding 2923 answers in total. Each question-answer pair was rephrased as a forecasting argument. For binary questions, we converted the pair into affirmative (or non-affirmative) statements of the forecasting question. For multiple-choice questions, we combined the answers with the question to form a forecasting argument. We generated these rephrasings using the Mistral-7B-Instruct-v0.3 LLM Jiang et al. (2023). All arguments were then manually reviewed, and any errors in the generated forecasting arguments were corrected. The questions and the generated claims are available in the supplementary material.
6.2.2 ArguCast with ArgLLMs
In order to populate ArguCast, we used LLMs to get predictions and ArgLLMs, leveraging LLMs, to construct QBAFs (we used the same prompts and hyperparameters as in Freedman et al. (2025)): we prompted the LLMs to generate supporting and attacking arguments; for each generated argument, we prompted the LLMs again to obtain an uncertainty score (acting as the base score). We investigated two variations of ArgLLMs: generating arguments at breadth $1,1$ (only generating a supporter and an attacker for each forecasting argument), and breadth $n,k$ (generating $n$ supporting and $k$ attacking arguments for each forecasting argument). We set the base score for each forecasting argument to 0.5. Figure 3 shows an example (at breadth $1,1$).
After obtaining the predictions for the forecasting arguments and the QBAFs, we apply coherence.
Figure 3: An example of ArgLLM (with Mistral-7B-Instruct-v0.3 as the underpinning LLM) applied to the forecasting argument $f$ at the top, giving a QBAF with a supporting and an attacking argument, with base scores 0.6 and 0.4 respectively, leading to a strength $\sigma(f)\!=\!0.6$ with DF-QuAD (given $\tau(f)\!=\!0.5$). The resulting forecaster QBAF leads to an incoherent forecast, as the 40% prediction is lower than $\xi_{2}(f)\!=\!0.5$, but $\sigma(f)\!>\!\xi_{1}(f)\!=\!0.5$.
6.3 Results
We compared the overall accuracy with the accuracy of the coherent predictions (i.e., if the coherent predictions have a higher accuracy, then coherence constraints would have a positive impact). As LLMs, we used the Mistral (Mistral-7B-Instruct-v0.3) Jiang et al. (2023), Mixtral (Mixtral-8x7B-Instruct-v0.1) Jiang et al. (2024), Llama 3 (Meta-Llama-3-8B-Instruct) Dubey et al. (2024), and GPT-4o (GPT-4o-mini) OpenAI (2024) models. Figure 3 illustrates an incoherent and incorrect prediction, which will be filtered out leading to a more accurate prediction.
Table 2 reports the forecasting accuracy for each LLM and the forecasting accuracy in the set of coherent predictions for both breadth $1,1$ and breadth $n,k$. The results indicate that applying coherence generally leads to improvements or, at least, preserves forecasting accuracy across all models. Note that applying coherence significantly reduces the total number of predictions, at breadth $n,k$ generally retaining more samples compared to breadth $n,k$. Most notably, Llama 3 retains only 44% of samples at breadth $1,1$, while it preserves 71% at breadth $n,k$. The most samples retained were with the Mistral model at breadth $n,k$, preserving 78% of the samples.
Raw LLM ArgLLM(B1,1) ArgLLM(Bn,k)
Total Acc. Total Coherent Acc. Total Coherent Acc.
Mistral 2923 77% 1621 80% 2280 81%
Mixtral 2923 75% 1693 78% 1668 75%
Llama 3 2923 75% 1294 82% 2064 79%
GPT-4o 2923 69% 1716 80% 1922 73%
Table 2: Accuracy comparison between raw LLMs and ArgLLMs with breadth $1,1$ and breadth $n,k$. For raw LLMs, we show accuracy as (correct/total). For ArgLLMs we show the percentages of coherent and correct predictions within those of coherent predictions (coherent & correct/total coherent). The bold percentage shows the best variant for the given model. The total columns show the number of samples retained for each model.
The key finding is that accuracy consistently improves when coherence is applied, particularly at breadth $1,1$. GPT-4o shows the most dramatic improvement, with accuracy increasing from 69% to 80% at breadth $1,1$. Llama 3 and Mixtral also achieve their best performance when coherence is applied at breadth $1,1$, reaching 82% and 78% accuracy, respectively. Interestingly, Mistral is the only model that achieved a higher accuracy at breadth $n,k$ compared to breadth $1,1$, which could suggest that already better performing models may require more arguments to be generated.
These results support a positive answer to RQ2. Notably, the improvements are most consistent at breadth $1,1$. Although small, the consistency of these improvements is valuable, as in the case of RQ1.
7 Crowd-sourced Experiments
Given the positive answers to RQ1 and RQ2, in this section we report on the findings from our last user study with participants drawn from a crowd-sourcing platform, which aimed to assess the prevalence of coherence by users in the real world. The participants were shown a number of forecasting debates following the ArguCast method, of different complexity. We then assessed participants’ conformance to our notion of coherence.
Figure 4: Example of a debate shown to users. This includes a forecasting argument (in cyan), a supporter thereof (in gray with a green solid arrow pointing towards the forecasting argument), an attacker of the forecasting argument (also in gray but with a red dotted arrow pointing towards the forecasting argument), and the votes of a fictitious forecaster $u$ (where blue thumbs up represents $V(u,a)=+$ and pink thumbs down represents $V(u,a)=$-).
7.1 Research Questions
We aim to evaluate the following RQs:
RQ3 Does our notion of coherence align with users’ reasoning in the context of forecasting?
RQ4 Does our notion of coherence align more with users’ reasoning when debates are not “complex”?
We anticipated, for RQ3, the null hypothesis (stating that users’ reasoning and our notion of coherence disagree in the same manner) not to be rejected. However, for RQ4, we expected the null hypothesis to be rejected (i.e., that RQ4 is answered affirmatively).
We focused on three “complexity” types, along three axes: vote, breadth, and depth given below (for more details see Appendix D).
Definition 3.
A simple variant is an ACF $\langle\mathcal{X}=\mathcal{F}\cup\mathcal{D},\mathcal{R}=\mathcal{A}\cup\mathcal{S},\mathcal{U},\mathcal{V},\mathcal{P}\rangle$ such that:
$|\mathcal{F}|=1\land|\mathcal{D}|=2\land|\mathcal{A}|=1\land|\mathcal{S}|=1\land(a,f)\in\mathcal{A}\rightarrow$
$\mathcal{V}(u,a)=+\land(s,f)\in\mathcal{S}\rightarrow\mathcal{V}(u,s)=-.$
So, a simple variant has three arguments: the forecasting argument, and a supporter and an attacker thereof; and the forecaster disagrees with the supporter and agrees with the attacker.
Definition 4.
A vote complex variant is an ACF $\langle\mathcal{X}=\mathcal{F}\cup\mathcal{D},\mathcal{R}=\mathcal{A}\cup\mathcal{S},\mathcal{U},\mathcal{V},\mathcal{P}\rangle$ such that:
$(\exists_{b\in\mathcal{D}}\exists_{f\in\mathcal{F}}(b,f)\in\mathcal{A}\land\mathcal{V}(u,b)=-)\lor$
$(\exists_{c\in\mathcal{D}}\exists_{d\in\mathcal{D}}\exists_{g\in\mathcal{F}}(c,g)\in\mathcal{A}\land(d,g)\in\mathcal{S}\land\mathcal{V}(u,c)=\mathcal{V}(u,d)).$
Thus, a vote complex variant is where the forecaster disagrees with an argument that attacks the forecasting argument (giving a double negative), or if the forecaster has conflicting votes (e.g. if the forecaster agrees with both arguments that support and attack the forecasting argument). The votes become more complex because the idea of disagreeing with an argument in an attack relation and conflicting votes is harder for some forecasters to understand [^10] We found this in a small pilot experiment with 12 participants..
Definition 5.
A breadth complex variant is an ACF $\langle\mathcal{X}=\mathcal{F}\cup\mathcal{D},\mathcal{R}=\mathcal{A}\cup\mathcal{S},\mathcal{U},\mathcal{V},\mathcal{P}\rangle$, such that:
$\lvert\{a\in\mathcal{D}|\not\exists_{r\in\mathcal{R}}{(r=(a_{1},a),a_{1}\in\mathcal{D})}\}\rvert=3$.
A breadth complex variant occurs when the forecaster adds a new supporter or attacker to the forecasting argument, i.e., the forecasting argument has three child arguments [^11] Note that since the forecaster adds the new argument, their vote on that argument is automatically set to agreement..
Definition 6.
A depth complex variant is an ACF $\langle\mathcal{X}=\mathcal{F}\cup\mathcal{D},\mathcal{R}=\mathcal{A}\cup\mathcal{S},\mathcal{U},\mathcal{V},\mathcal{P}\rangle$ such that:
$|\{a\in\mathcal{D}|\exists_{r\in\mathcal{R}}{(r=(a,a_{1}),a_{1}\in\mathcal{X})}\}|=1$.
A depth complex variant occurs when the forecaster adds a new argument to any of the leaf (i.e. unattacked and unsupported) nodes. This argument can either support or attack one of the child arguments (the supporter or attacker of the forecasting argument).11
With these notions of complexity, we can refine RQ4 into:
RQ4.1 Does our notion of coherence align more with users’ reasoning when debates are not vote complex?
RQ4.2 Does our notion of coherence align more with users’ reasoning when debates are not breadth complex?
RQ4.3 Does our notion of coherence align more with users’ reasoning when debates are not depth complex?
7.2 Experimental Design
In order to answer the RQs we ran a user experiment on the online recruitment platform Prolific [^12] Prolific (www.prolific.com) was chosen for its superior data quality compared to other similar platforms Péer et al. (2021). using a Qualtrics survey. In the survey, each user was assigned four variants of a debate, chosen randomly between two available debates for two different binary forecasting questions, with abstract entities to prevent bias (i.e., Will X win their next tennis match against Y? and Will Party A win the next election against Party B?), where each debate has 60 potential variants. The (variants of the) debates capture votes of a fictitious forecaster Alex and Alex’s prediction on the forecasting question. The full debates are shown in Appendix C. Figure 4 gives a simple example thereof.
The survey follows this design flow: (1) Main Test (4 x 3 questions) – here, we provide the users with four (variants of) debates, each containing Alex’s prediction, part of the debate with Alex’s votes, and three questions, consisting of an attention check, whether the user agrees with Alex’s prediction, and their feedback on their reasoning (optional). This task addresses both research questions (RQ3 and RQ4). (2) Post-tasks (4 questions) – the participants are asked questions regarding their experience with the questionnaire.
In this experiment, we set $\xi_{1}(f)=\xi_{2}(f)=0.5$. We choose $\xi_{2}(f)=0.5$ due to the forecasting questions being fictitious and having a 50% probability of occurring. The participants were divided into two groups: 50% of them were asked to assess simpler questions first and more complex questions later (named the Simple-Complex group) while the other 50% of were asked to assess more complex questions first and simpler questions after (the Complex-Simple group). In both cases, users were randomly allocated (50%, 50%) which (of the two) forecasting questions to assess. For the four (variants of) debates the users saw, Alex’s forecast was also allocated randomly (1/3, 1/3, 1/3) amongst $50\%,=50\%,>50\%$. In Appendix E we further discuss the experimental design of this study.
For an expected significance of 0.05, a power of 0.8, and an effect size of 0.3, we needed 174 samples for RQ3 (where a $\chi^{2}$ test is used) and 90 for each RQ4.i. We ended up recruiting 100 participants as this yielded 400 samples for RQ3 and 200 samples for each RQ4.i, or 360 for RQ3 and 180 for each RQ4.i accounting for a 10% dropout, which was sufficient.
7.3 Participant Recruitment and Ethics Statement
We recruited 100 anonymous English-speaking participants aged 18+ from Prolific, paying them an average of £4.50 for their time, based on the platform’s standard reimbursement rate of £26.82 per hour. The data collection was anonymous. Participants were asked to provide informed consent after reading the participant information sheet, and they could withdraw at any time before completing the survey by simply closing the questionnaire window, with their data permanently deleted. Participants were informed that once they completed the survey, they could no longer withdraw, and we would retain their data. Ethics approval was obtained from our institution prior to conducting the experiments.
7.4 Results
Out of 100 participants, only 46 passed all attention checks and were included in our study. For RQ3 we analysed 184 samples using McNemar’s test McNemar (1947) used for determining whether two classification models disagree in the same way. The null hypothesis was that the two marginal probabilities for each outcome (from our notion and users) were identical (i.e., H0: $12=76$, H1: $12\neq 76$ from the contingency Table 3). The chi-squared test (commonly used in conjunction with McNemar’s) with one degree of freedom yielded $p<.00001$. Given the result is significant at $p<0.05$, we rejected H0 in favour of H1, indicating that the users’ opinion of coherence does not align with our notion.
User Coherent?
Yes No
Model Coherent? Yes 44 12
No 76 52
Table 3: Contingency table for McNemar’s test. Columns/rows represent the number of debate variants the users/our notion of coherence (respectively) found coherent for the same debate variants.
s v b d d/b v/d v/b v/d/b
Aligned 20 7 7 9 9 8 9 27
[…]
Table 4: Number of (not) aligned and pairs (alignment between our notion of coherence and the participants in the experiment) for the debate variations: simple, breadth complex, depth complex and vote complex.
Table 5 summarises the results for RQ4.1-RQ4.3. For each complexity type, we conducted a two-sample one sided t-test to determine if complex debate variants aligned more with our notion of coherence compared to the simpler variants. Our null hypothesis for all RQ4.1-RQ4.3 was H0: $\mu_{\text{complex}}\leq\mu_{\neg\text{complex}}$ against the alternative H1: $\mu_{\text{complex}}>\mu_{\neg\text{complex}}$. All three complexity types were statistically significant at $p<0.05$. Vote complexity yielded a p-value of 0.009, breadth complexity a p-value of 0.04, and depth complexity the most significant effect with a p-value of 0.002. Based on these results, we can reject all three null hypotheses.
Complexity $\mu_{\text{complex}}$ $\mu_{\neg\text{complex}}$ $SD_{\text{complex}}$ $SD_{\neg\text{complex}}$ p-value
Vote 0.57 0.48 0.245 0.25 0.009
Breadth 0.55 0.49 0.247 0.25 0.04
Depth 0.58 0.47 0.24 0.25 0.002
Table 5: Mean and standard deviation of aligned pairs (between our notion of coherence and the participants from the user experiment) for each complex and not-complex debate variations. These values are derived from Table 4. Then, it reports on the p-value derived from the two-sample one sided t-test.
Rejecting RQ3 suggests that the notion of coherence with $\xi_{1}(f)=\xi_{2}(f)=0.5$ does not accurately reflect the intuition of what users perceive as coherent. This further implies the need for a coherence constraint, as users are not naturally coherent in their judgment. Additionally, since we have rejected RQ4.1, RQ4.2, and RQ4.3 and these were refinements of RQ4, we can also reject RQ4. This means that our notion of coherence aligns more with users’ reasoning when debates are complex. This could suggest that our notion of coherence is more intuitive when more information is available to the users. Furthermore, this indicates that with complex debate variations, the users spend more cognitive effort with complex debate variations, which helps them counteract their biases. We have left the qualitative error analysis of forecaster biases to future work.
8 Conclusion
In this paper, we introduced a novel notion of argumentative coherence for judgmental forecasting, formally defined in terms of alignment between forecasters’ internal argumentative reasoning and their predictions. Then, through crowd-sourced and in-person user studies, as well as experiments with LLMs, we systematically evaluated the usefulness of the coherence property.
First, in an in-person user study with ArguCast, we found that filtering out incoherent forecasts improved the aggregated predictions, moving them closer to expert baselines. Second, in the experiments with LLMs, we applied coherence to LLM-generated forecasts and ArgLLM-generated argumentative reasoning and demonstrated consistent improvements in accuracy across different models. These results from the in-person and LLM experiments support the practical need for integrating argumentative coherence into judgmental forecasting systems (ArguCast, GJOpen, Metaculus). Finally, in the crowd-sourced experiments, we found that the users’ opinions may not align with (our notion of) coherence. This suggests that some form of coherence constraints are necessary. Additionally, we observe that the users’ opinions aligned more with our notion of coherence when debates were more complex. We speculate that this is due to our notion being more intuitive when more information is available. The results also indicated that as users spend more cognitive effort with complex debates (which implies that they reason more), it helps them counteract their biases. This may have implications for works considering argumentative models of reasoning, e.g. Rago et al. (2023).
For future work, we plan to run a similar study to the crowd-sourced study with argumentation experts to determine whether our notion of coherence is intuitive to them. Additionally, we aim to extend our notion of coherence to provide a degree of coherence rather than a binary outcome. This would allow us to weight predictions, giving more influence to those that are more coherent, thereby enhancing the impact on the group forecast. We also plan to explore relationships between our approach and probabilistic argumentation Toni et al. (2023). Finally, we would like to use FORECAST2023 Gorska et al. (2024) and/or the hybrid-forecasting competition Benjamin et al. (2023) dataset to populate an argumentation-based forecasting system (e.g. ArguCast) and study whether the notion of coherence improves forecasting accuracy, as was the case for the notion of coherence in Irwin et al. (2022).
Acknowledgments
This research was partially funded by the ERC under the EU’s Horizon 2020 research and innovation programme (grant agreement no. 101020934, ADIX), by J.P. Morgan and by the Royal Academy of Engineering, UK, under the Research Chairs and Senior Research Fellowships scheme (grant agreement no. RCSRF2021\11\45).
Supplementary Material
Appendix A Desirable Properties of Forecaster QBAFs
We introduce two properties that provide sanity checks for the forecaster QBAFs, which is the underlying framework for our notion of coherence.
Property 1.
(Neutral Forecasting) $\forall f\in\mathcal{F}$:
$\forall_{a\in\mathcal{D}}{\mathcal{V}(u,x)\in\{?,\text{\small undefined}\}}\rightarrow\sigma(f)=0.5$
This property ensures that if a forecaster has not voted on any argument then the strength of the forecasting argument $f$ should be $0.5$, which is the base score of the forecasting argument. So, based on Definition 2, if the forecaster has not expressed their opinion then their prediction should be close to $\xi_{2}(f)$. This will force forecasters to vote (i.e., show their reasoning) if they want to make a prediction on $f$ further away from $\xi_{2}(f)$.
Proposition 1.
Forecaster QBAFs satisfy the Neutral Forecasting property.
Proof.
We need to show that in forecaster QBAFs, if $\mathcal{V}(a)=?$ $\forall a\in\mathcal{X}$ then $\forall f\in\mathcal{F}$ $\sigma(f)=0.5$ holds. It can be seen, by Definition 1, that if $a\not\in\mathcal{F}$ and $\mathcal{V}_{u}(a)=?$ then $\tau(a)=0$, and if $a\in\mathcal{F}$ then $\tau(a)=0.5$. Since the forecaster QBAF is a tree, we first consider some $a$ which is a leaf, i.e., $\nexists b\in\mathcal{X}$ such that $(b,a)\in\mathcal{A}\cup\mathcal{S}$, by the definition of DF-QuAD, we know that $\sigma(a)=\tau(a)$. Thus, for any $f\in\mathcal{F}$ which is a leaf, the property holds. If $\exists f\in\mathcal{F}$ which are not leaves, we calculate the strengths of the arguments which are only attacked or supported by leaves, and as shown by the fact that DF-QuAD satisfies Amgoud and Ben-Naim [2018] Principle 6 (neutrality), as well as the fact that, by Definition 1, $(\mathcal{F}\times\mathcal{X})\cap(\mathcal{A}\cup\mathcal{S})=\emptyset$, we know that for these arguments also, $\sigma(a)=\tau(a)$ holds by the definition of DF-QuAD. We then continue to propagate up the tree in the same way, until $\forall f\in\mathcal{F}$, it is shown that $\sigma(f)=0.5$, and thus the property holds.
Note that this is not the case from existing work Gorur et al. [2023], consider a single attacker $(a,f)\in\mathcal{A}^{u}$ of the forecasting argument with $V_{u}(a)=?$, then the strength for the forecasting argument would be $\sigma(f)=0.25$. This is a counterexample that shows the Neutral Forecasting property is not satisfied for their notion of coherence.
Property 2.
(Bias-based Forecasting) For $f\in\mathcal{F}$:
$\exists_{(a,f)\in\mathcal{A}_{u}}\tau_{u}(a)=1\land\forall_{x\in\mathcal{D}\setminus\{a\}}\mathcal{V}(u,x)\in\{?,\text{\small undefined}\}$
$\rightarrow\sigma(f)<0.5\ \text{and}$
$\exists_{(s,f)\in\mathcal{S}_{u}}\tau_{u}(s)=1\land\forall_{x\in\mathcal{D}\setminus\{s\}}\mathcal{V}(u,x)\in\{?,\text{\small undefined}\}$
$\rightarrow\sigma(f)>0.5$
If the forecaster agrees with only the attackers (or supporters), the final strength $\sigma(f)$ should be less (or more) than $0.5$. This differs from the notion in Gorur et al. [2023], where when there were many (more than 50) supporters (or attackers) which the forecaster had no opinion and the only argument with which the forecaster agrees is an attacker (or supporter), then the strength would be exactly $\sigma(f)=0.5$, which is the base score. This would identify forecasters who exhibit bias, as their prediction would deviate from their opinions.
Proposition 2.
Forecaster QBAFs satisfy the Bias-based Forecasting property.
Proof.
Consider the first case, in which we need to show that $\sigma(f)<0.5$. The proof showing that $\forall x\in\mathcal{X}\setminus\{f\}$, $\sigma(x)=\tau(x)$ is the same as that for Proposition 1. We thus know that $\sigma_{\mathcal{S}}^{u}(f)=0$ and then, since $\sigma(a)=\tau(a)=1$, $\sigma_{\mathcal{A}}^{u}(f)=1$. Thus, it can be seen by the definition of DF-QuAD that $\sigma(f)<\tau(f)=0.5$ if $\sigma_{\mathcal{A}}^{u}(f)>\sigma_{\mathcal{S}}^{u}(f)$ (this was shown in Potyka [2021] (strict monotonicity) excluding when $\tau(f)=0$ and $\tau(f)=1$, which is not the case here).
The proof for the second case is analogous, therefore, the Bias-based Forecasting property is satisfied for forecaster QBAFs as both conditions of the property are satisfied.
So, our notion of coherence has advantages over that in Gorur et al. [2023], which satisfies neither of these properties. We leave a full theoretical analysis of our method’s properties to future work, given the focus on the user studies in this paper.
Appendix B Debates Generated in the In-Person Experiments
The debates generated for the three questions in the in-person experiment are given in Figure 5, 6, and 7.
Figure 5: Supporting and attacking arguments generated for the Euro question by the participants of the in-person experiment.
Figure 6: Supporting and attacking arguments generated for the Labour question by the participants of the in-person experiment.
Figure 7: Supporting and attacking arguments generated for the Nvidia question by the participants of the in-person experiment.
Appendix C Full Debates on Forecasting Questions
The full debates used in our experiments are given in Figure 8 and Figure 9.
Figure 8: Full debate for the forecasting question ‘Will X win their next tennis match against Y?’.
Figure 9: Full debate for the forecasting question ‘Will Party A win the next election against Party B?’.
Appendix D Abstract Structure of Complexities
In this section we represent the abstract structure of all debate variation complexities (see Figure 10, 11, 12, 13, 12, 14, 15, 16, and 17).
Figure 10: Abstract structure of the simple debate variant.
Figure 11: Abstract structure of the vote complex debate variant.
Figure 12: Abstract structure of the breadth complex debate variant.
Figure 13: Abstract structure of the depth complex debate variant.
Figure 14: Abstract structure of the vote/breadth complex debate variant.
Figure 15: Abstract structure of the vote/depth complex debate variant.
Figure 16: Abstract structure of the depth/breadth complex debate variant.
Figure 17: Abstract structure of the vote/depth/breadth complex debate variant.
Appendix E Experimental Design
The participants will be split such that 50% of them assess simpler questions first and more complex questions after (Simple-Complex) and 50% of them assess more complex questions first and simpler questions after (Complex-Simple). In both cases, users will be randomly allocated (50%, 50%) which forecasting question they will assess (amongst Q1,Q2). For the 4 (variants of) debates the users will see, Alex’s forecast is also allocated randomly (1/3, 1/3, 1/3) amongst $50\%,=50\%,>50\%$.
  • For the Simple-Complex case, they will be first shown the simple debate variant. We will then split the participants such that 1/3 assess vote complex debate variants, 1/3 assess breadth complex debate variants, and 1/3 assess depth complex debate variants.
    • From vote complex debate variants participants will be split into 50% vote/depth complex debate variants or 50% vote/breadth complex debate variants.
    • From breadth complex debate variants participants will be split 50% vote/breadth complex debate variants or 50% depth/breadth complex debate variants.
    • From depth complex debate variants participants will be split 50% vote/depth complex debate variants or 50% depth/breadth complex debate variants.
        Then, regardless of what type of debate variants they get, they will be shown a vote/depth/breadth complex debate variant.
  • For the Complex-Simple case they will be shown the vote/depth/breadth complex debate variants. We will then split further the participants such that 1/3 assess vote/depth complex debate variants, 1/3 assess vote/breadth complex debate variants, and 1/3 assess depth/breadth complex debate variants.
    • From vote/depth complex debate variants, participants will be split into 50% vote complex debate variants or 50% depth complex debate variants.
    • From vote/breadth complex debate variants, participants will be split 50% vote complex debate variants or 50% breadth complex debate variants.
    • From depth/breadth complex debate variants, participants will be split into 50% depth complex debate variants or 50% breadth complex debate variants.
        Then, regardless of what type of complex debate variant they get, they will be shown the simple variant.
Figure 18: Experimental design for our user experiments where subfigure a) shows the experimental design for the Simple-Complex group and subfigure b) shows the experimental design for the Complex-Simple group.